# Models of Set Theory: Extensions and Dead-ends

### v5

@@ 2. PRELIMINARIES
(b) $\mathcal{M}$ carries a definable global well-ordering, i.e., for some $p\in M$ and some set-theoretic formula $\varphi(x,y,z)$, $\mathcal{M}$ satisfies $\varphi(x,y,p)$ well-orders the universe”.
@@
(c) $\mathcal{M}$ carries a definable global choice function, i.e., for some $p\in M$ and some set-theoretic formula $\psi(x,y,z)$, $\mathcal{M}$ satisfies $\psi(x,y,p)$ is the graph of a global choice function”.
@@ 3. TALLER CONSERVATIVE ELEMENTARY EXTENSIONS
3.1. Theorem. Every model $\mathcal{M}$ of $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ (of any cardinality) has a conservative elementary extension $\mathcal{N}$ such that $\mathcal{N}$ is taller than $\mathcal{M}$.
We shall present two proofs of Theorem 3.1, the first is based on a class-sized syntactic construction taking place within a model of set theory; it was inspired by Kaufmann’s proof of the MacDowell-Specker theorem using the Arithmetized Completeness Theorem, as described by Schmerl [Schm-1]. The second one is based on a model-theoretic argument that is reminiscent of the ultrapower proof of the McDowell-Specker theorem (as in [En-5]). The first proof is short and devilish; the second proof is a bit longer but is more transparent due to its combinatorial flavor; it also lends itself to fine-tuning, as indicated in Remark 3.3 and Theorem 3.4.
@@
[…]
@@
Starting with a model $\mathcal{M}$ of $\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right),$ by Theorem 2.14, $\mathcal{M}$ carries a global definable well-ordering. We will construct an increasing sequence of $\mathcal{L}_{\mathrm{set}}$-structures $\left\langle\mathcal{N}_{k}:k\in\omega\right\rangle$ that satisfies the following properties for each $k\in\omega$:
(1) $\ \ \mathcal{N}_{0}=\mathcal{M}.$
(2) $\ \ \mathcal{M}\preceq_{\Pi_{k+2}}\mathcal{N}_{k}\preceq_{\Pi_{k+1}}\mathcal{N}_{k+1}$.
(3)   There is an ordinal in $\mathcal{N}_{1}$ that is above all of the ordinals of $\mathcal{M}$.
(4) $\ \ \mathcal{N}_{k}$ is a conservative extension of $\mathcal{M}$.
Thus
$\mathcal{M=N}_{0}\preceq_{\Pi_{1},\ \mathrm{cons,\ taller}}\mathcal{N}_{1}\preceq_{\Pi_{2},\ \mathrm{cons}}\mathcal{N}_{2}\preceq_{\Pi_{3},\ \mathrm{cons}}\mathcal{N}_{3,\ \mathrm{cons}}\cdot\cdot\cdot$
and for each $k\in\omega$, $\mathcal{M}\preceq_{\Pi_{n+2},\ \mathrm{cons}}\mathcal{N}_{k}.$ This shows that with the choice of $\mathcal{N}:=\bigcup\limits_{n\in\omega}\mathcal{N}_{n}$, we have $\mathcal{M}\prec_{\mathrm{cons}}\mathcal{N}$; and $\mathcal{N}$ is taller than $\mathcal{M}$. So the proof will be complete once we explain how to recursively build the desired chain $\left\langle\mathcal{N}_{k}:k\in\omega\right\rangle.$
Using the notation of Definition 2.7, let $\mathcal{L}_{\mathrm{set}}^{+}=$ $\mathcal{L}_{\mathrm{set}}\cup\{c_{a}:a\in\mathrm{V}\}$ be the language $\mathcal{L}_{\mathrm{set}}$ of set theory with a constant $c_{a}$ for each $a\in\mathrm{V}$. To build $\mathcal{N}_{1}$ we argue within $\mathcal{M}$: add a new constant $c$ to $\mathcal{L}_{\mathrm{set}}^{+}$ and consider the theory $T_{1}$ defined as follows:
$T_{1}=\left\{\mathrm{Ord}(c)\right\}\cup\{c_{\alpha}\in c:\mathrm{Ord}(\alpha)\}\cup\mathrm{Th}_{\Pi_{3}}($$V$$,\in,c_{a})_{a\in\mathrm{V}}.$
Thus $T_{1}$ is a proper class within $\mathcal{M}$. Arguing within $\mathcal{M}$, and using Fact 2, it is easy to see that $\mathrm{Con}(T_{1})$ holds, and therefore by Fact 1 we can get hold of a proper class model $\mathcal{N}_{1}$ of $T_{1}$. Clearly $\mathcal{M}\prec_{{}_{\Pi_{3}}}\mathcal{N}_{1}$, there is an ordinal in $\mathcal{N}_{1}$ that is above all of the ordinals of $\mathcal{M}$, and $\mathcal{N}_{1}$ is a conservative extension of $\mathcal{M}$ (by the last clause of Fact 1).
@@
$T_{k}=\mathrm{Th}_{\Pi_{k+1}}(\mathcal{N}_{k},c_{b})_{b\in N_{k}}\cup\mathrm{Th}_{\Pi_{k+2}}($$V$$,\in,c_{a})_{a\in\mathrm{V}}.$
@@
Fact 2 together with the inductive assumption that $\mathcal{M}\prec_{\Pi_{k+1}}\mathcal{N}_{k}$ makes it evident that $\mathrm{Con}(T_{k})$ holds in $\mathcal{M}$. Hence by Fact 1 we can get hold of a model $\mathcal{N}_{k+1}$ of $T$ whose elementary diagram is definable in $\mathcal{M}$. This makes it clear that $\left\langle\mathcal{N}_{i}:1\leq i\leq k+1\right\rangle$ meets (1) through (4). $\square$
Second proof of Theorem 3.1. Given $\mathcal{M}\models\mathrm{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ we will construct an ultrafilter $\mathcal{U}$ over the $\mathcal{M}$-definable subsets of $\mathrm{Ord}^{\mathcal{M}}$ satisfying the following properties:
(1)   Every member of $\mathcal{U}$ is unbounded in $\mathrm{Ord}^{\mathcal{M}}$.
(2)   Given any $\mathcal{M}$-definable $f:\mathrm{Ord}^{\mathcal{M}}\rightarrow\mathcal{M}$, there is a class $X\in\mathcal{U}$ such that either $f\upharpoonright X$ is one-to-one, or the range of $f\upharpoonright X$ is a set (as opposed to a class).
(3)   Given any $n\in\mathbb{\omega}$, the family $\mathcal{U}_{n}=\{X\in\mathcal{U}:X$ is $\Sigma_{n}$-definable in $\mathcal{M}$} is $\mathcal{M}$-definable in the sense that there is a parametric formula $\psi_{n}(\alpha,\beta)$ such that:
$X\in\mathcal{U}_{n}$ iff $\exists\alpha\in\mathrm{Ord}^{\mathcal{M}}$ $X=\{\beta\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models\psi_{n}(\alpha,\beta)\}.$
We will build $\mathcal{U}$ in $\omega\times\mathrm{Ord}^{\mathcal{M}}$ stages in the following sense: for each $n\in\omega$ and $\alpha\in\mathrm{Ord}^{\mathcal{M}}$ we shall employ external recursion on $n$ and an internal recursion on $\alpha$ to a parametrically $\mathcal{M}$-definable subset $X_{n,\alpha}$ of $\mathrm{Ord}^{\mathcal{M}}$, and then we can define $\mathcal{U}$ as:
$\mathcal{U}=\{X_{n,\alpha}:n\in\omega,\alpha\in\mathrm{Ord}^{\mathcal{M}}\}\cup\mathcal{U}_{-1},$
where $\mathcal{U}_{-1}$ denote the collection of co-bounded parametrically $\mathcal{M}$-definable subsets of $\mathrm{Ord}^{\mathcal{M}}$, i.e., parametrically $\mathcal{M}$-definable subsets $X$ of $\mathrm{Ord}^{\mathcal{M}}$ such that, as viewed in $\mathcal{M}$, the complement of $X$ is a set (as opposed to a proper class). Let $\left\langle f_{n,,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ denote an enumeration within $\mathcal{M}$ of all $\Sigma_{n}$-functions of $\mathcal{M}$ with domain $\mathrm{Ord}^{\mathcal{M}}$. Note that such an $\mathcal{M}$-enumeration exists because:
@@
(a) By Theorem 2.9 for each $n\in\omega$ there is a definable satisfaction predicate for $\Sigma_{n}$-formulae, and
@@
(b) By Theorem 2.14 $\mathcal{M}$ has a definable global well-ordering.
Suppose for some $n\in\omega$ we have constructed $\mathcal{U}_{n}=\{X_{n,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\}$ satisfying the following inductive clauses:
$(i)$ $\ \ \mathcal{U}_{n}$ is a parametrically $\mathcal{M}$-definable family of classes in $\mathcal{M}$, i.e., there is a formula $\psi_{n}(\alpha,\beta)$ (with suppressed parameters) such that:
$X\in\mathcal{U}_{n}$ iff $\exists\alpha$ $X=\{\beta\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models\psi_{n}(\alpha,\beta)\}.$
$(ii)$ $\ \ \mathcal{U}_{n}$ has the finite intersection property in the sense of $\mathcal{M}\mathfrak{,}$ i.e., if $\psi_{n}(\alpha,\beta)$ is as above, then $\mathcal{M}$ satisfies the sentence below:
@@
For all finite sequences of ordinals $\left\langle\alpha_{i}:i<k\right\rangle$ there is $\beta\in\mathrm{Ord}$ such that $\forall i<k\ \psi_{n}(\alpha_{i},\beta)$.
$(iii)$ $\ \ f_{n,\alpha}\upharpoonright X_{n,\alpha}$ is either one-to-one, or the range of $f_{n,\alpha}\upharpoonright X_{n,\alpha}$ is a set.
$(iv)$ $\ \ \mathcal{U}_{-1}\subseteq\mathcal{U}_{n}.$
We shall now describe a definable sequence $\left\langle X_{n+1,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ which continues to satisfy the clauses $(i)$ through $(iii)$ above, and we define $\mathcal{U}_{n+1}:=\left\{X_{n+1,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\}.$ Since this part of the construction involves an internal recursion within $\mathcal{M}$ we will argue as if we are inside $\mathcal{M}$ and therefore omit reference to $\mathcal{M}$ itself. Suppose we have built $\left\langle X_{n+1,\beta}:\beta<\alpha\right\rangle$ and let $\mathcal{F}_{n+1,\alpha}$ denote:
$\{X_{m,\delta}:m<n+1,\delta<\alpha\},$
and let $\mathcal{F}_{n+1,\alpha}^{\ast}$ denote the set of all finite intersections of $\mathcal{F}_{n+1,\alpha}$. Also, let $\left\langle Y_{\alpha}:\alpha\in\mathrm{Ord}\right\rangle$ be an enumeration of $\mathcal{F}_{n+1,\alpha}^{\ast}$. We now wish to define a class $X_{n+1,\alpha}$ which “takes care of” $f_{n+1,\alpha}$, and which meets every $Y_{\alpha}$ We distinguish the following two cases:
Case 1: There is an ordinal $\gamma$ such that for every ordinal $\alpha$ we have:
$f_{n+1,\alpha}^{-1}\{\beta:\beta<\gamma\}\cap Y_{\alpha}\neq\varnothing.$

### v6

@@ 2.14. Myhill-Scott Theorem [MS]The following statements are equivalent for $\mathcal{M}\models\mathsf{ZF}$:
    $\mathcal{M}$ carries a definable global well-ordering, i.e., for some $p\in M$ and some set-theoretic formula $\varphi(x,y,z)$, $\mathcal{M}$ satisfies $\varphi(x,y,p)$ well-orders the universe”.
  3. (c)
@@
    $\mathcal{M}$ carries a definable global choice function, i.e., for some $p\in M$ and some set-theoretic formula $\psi(x,y,z)$, $\mathcal{M}$ satisfies $\psi(x,y,p)$ is the graph of a global choice function”.
@@ 3. TALLER CONSERVATIVE ELEMENTARY EXTENSIONS
3.1. Theorem. Every model $\mathcal{M}$ of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right)$ (of any cardinality) has a conservative elementary extension $\mathcal{N}$ such that $\mathcal{N}$ is taller than $\mathcal{M}$.
We shall present two proofs of Theorem 3.1, the first one is based on a class-sized syntactic construction taking place within a model of set theory; it was inspired by Kaufmann’s proof of the MacDowell-Specker theorem using the Arithmetized Completeness Theorem, as described by Schmerl [Schm-1]. [^17] As noted by Phillips [P] and Gaifman [G-2], the McDowell-Specker proof lends itself to a fine-tuning that ensures that the conservative elementary extension $\mathcal{N}$ of a given model $\mathcal{M}$ has the minimality property, i.e., there is no $\mathcal{K}$ such that $\mathcal{M}\prec\mathcal{K}\prec\mathcal{N}$. However, Kaufmann’s slick proof does not seem to lend itself to this embellishment. The second one is based on a model-theoretic construction that is reminiscent of the original ultrapower proof [^18] The original proof by MacDowell and Specker of their result uses the concept of a “finitely additive 2-valued measure”, which is an alternative formulation of the notion of an ultrafilter. of the MacDowell-Specker theorem [MS]. The first proof is short and devilish; the second proof is a bit longer but is more transparent due to its combinatorial flavor; it also lends itself to a refinement, as indicated in Remark 3.3 and Theorem 3.4.
@@
FACT 1: Suppose $\mathcal{M}$ is a model of $\mathsf{ZF}$ that carries an $\mathcal{M}$-definable global well-ordering, and $T$ is an $\mathcal{M}$-definable class of first order sentences such that $\mathcal{M}$ satisfiesTT is a consistent first order theory”. Then there is a model $\mathcal{N}\models T^{\mathrm{st}}$ such that the elementary diagram of $\mathcal{N}$ is $\mathcal{M}$-definable, where $T^{\mathrm{st}}$ is the collection of sentences in $T$ with standard shape [^19] More specifically, $\varphi\in T^{\mathrm{st}}$ iff $\varphi\in T$ and there is some standard formula $\varphi^{\ast}$ such that $\mathcal{M}$ believes that $\varphi$ is the result of substituting constants for the free variables of $\varphi^{\ast}.$.
@@
Starting with a model $\mathcal{M}$ of $\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right),$ by Theorem 2.14, $\mathcal{M}$ carries a global definable well-ordering. We will construct an increasing sequence of $\mathcal{L}_{\mathrm{set}}$-structures $\left\langle\mathcal{N}_{k}:k\in\omega\right\rangle$ that satisfies the following properties for each $k\in\omega$:
  1. (1)
    $\mathcal{N}_{0}=\mathcal{M},$
  2. (2)
    $\mathcal{M}\preceq_{\Pi_{k+2}}\mathcal{N}_{k}\preceq_{\Pi_{k+1}}\mathcal{N}_{k+1}$.
  3. (3)
    There is some $\alpha\in\mathrm{Ord}^{\mathcal{N}_{1}}$ that is above each $\beta\in\mathrm{Ord}^{\mathcal{M}}$, thus $\mathcal{M}\preceq_{\mathrm{taller}}\mathcal{N}_{k}$ for all $k\geq 1$.
  4. (4)
    $\mathcal{N}_{k}$ is a self-extension of $\mathcal{M}$, and in particular $\mathcal{M}\preceq_{\mathrm{cons}}\mathcal{N}_{k}$. [^20] See Definition 2.4(b) and Remark 2.5(a).
In other words:
$\mathcal{M=N}_{0}\preceq_{\Pi_{3},\mathrm{taller}}\mathcal{N}_{1}\preceq_{\Pi_{2}}\mathcal{N}_{2}\preceq_{\Pi_{3}}\mathcal{N}_{3}\$...,
and for each $k\in\omega$, $\mathcal{M}\preceq_{\Pi_{k+2},\mathrm{cons,taller}}\mathcal{N}_{k}.$
In light of part (a) of Remark 2.5, this shows that $\mathcal{M}\prec_{\mathrm{cons,taller}}\mathcal{N}$, where $\mathcal{N}:=\bigcup\limits_{k\in\omega}\mathcal{N}_{k}$. So the proof will be complete once we explain how to recursively build the desired chain $\left\langle\mathcal{N}_{k}:k\in\omega\right\rangle.$
Using the notation of Definition 2.7, let $\mathcal{L}_{\mathrm{set}}^{+}=$ $\mathcal{L}_{\mathrm{set}}\cup\{\dot{a}:a\in\mathrm{V}\}$ be the language $\mathcal{L}_{\mathrm{set}}$ of set theory with a constant $\dot{a}$ for each $a\in\mathrm{V}$. To build $\mathcal{N}_{1}$ we argue within $\mathcal{M}$: add a new constant $c$ to $\mathcal{L}_{\mathrm{set}}^{+}$ and consider the theory $T_{1}$ defined as follows:
$T_{1}=\left\{\mathrm{Ord}(c)\right\}\cup\{\dot{\alpha}\in c:\mathrm{Ord}(\alpha)\}\cup\mathrm{Th}_{\Pi_{3}}($$V.,\in,\dot{a})_{a\in\mathrm{V}}$.
Thus $T_{1}$ is a proper class within $\mathcal{M}$. Arguing within $\mathcal{M}$, and using Fact 2, it is easy to see that $\mathrm{Con}(T_{1})$ holds, and therefore by Fact 1 we can get hold of a proper class model $\mathcal{N}_{1}$ of $T_{1}$ such that $\mathcal{N}_{1}$ is a self-extension of $\mathcal{M}$. Thus $\mathcal{M}\prec_{{}_{\Pi_{3,\mathrm{cons}}}}\mathcal{N}_{1}$, and there is an ordinal in $\mathcal{N}_{1}$ that is above all of the ordinals of $\mathcal{M}$.
@@
$T_{k+1}=\mathrm{Th}_{\Pi_{k+1}}(\mathcal{N}_{k},b)_{b\in N_{k}}\cup\mathrm{Th}_{\Pi_{k+2}}($$V.,\in,a)_{a\in\mathrm{V}}$.
@@
Fact 2 together with the inductive assumption that $\mathcal{M}\prec_{\Pi_{k+1}}\mathcal{N}_{k}$ makes it evident that $\mathrm{Con}(T_{k+1})$ holds in $\mathcal{M}$. Hence by Fact 1 we can get hold of a model $\mathcal{N}_{k+1}$ of $T$ such that $\mathcal{N}_{k+1}$ is a self-extension of $\mathcal{M}$. This makes it clear that $\left\langle\mathcal{N}_{i}:1\leq i\leq k+1\right\rangle$ meets (1) through (4). $\square$
Second proof of Theorem 3.1. Given $\mathcal{M}\models\mathsf{ZF}+\exists p\left(\mathrm{V}=\mathrm{HOD}(p)\right),$ let $\mathbb{B}$ be the Boolean algebra consisting of $\mathcal{M}$-definable subsets of $\mathrm{Ord}^{\mathcal{M}}$ (ordered by $\subseteq$). In Stage 1 of the proof we will build an appropriate ultrafilter $\mathcal{U}$ on $\mathbb{B},$ and in Stage 2 we will verify that the definable ultrapower of $\mathcal{M}$ modulo $\mathcal{U}$, denoted $\mathcal{M}^{\mathrm{Ord}^{\mathcal{M}}}/\ \mathcal{U}$, is a conservative elementary extension of $\mathcal{M}$ that is taller than $\mathcal{M}$.
STAGE 1: The following facts come handy in this stage:
@@
Fact 1. By Theorem 2.9, for each $n\in\omega$ there is a definable satisfaction predicate for $\Sigma_{n}$-formulae.
@@
Fact 2. By Theorem 2.14, $\mathcal{M}$ has a parametrically definable global well-ordering.
To facilitate the construction of $\mathcal{U}$, let us introduce some conventions:
  1. $(i)$
    Given $\mathfrak{X}\subseteq\mathcal{P}(M)$, we say that $\mathfrak{X}$ is $\mathcal{M}$-listable, if that there is a parametric formula $\psi(\alpha,\beta)$ such that for all $X\subseteq M$ we have:
$X\in\mathfrak{X}$ iff $\exists\alpha\in\mathrm{Ord}^{\mathcal{M}}$ $X=\{m\in M:\mathcal{M}\models\psi(\alpha,m)\}.$
  1. In the above context, if $\psi$ is a $\Sigma_{n}$-formula, $\mathfrak{X}$ is said to be $\Sigma_{n}$-listable in $\mathcal{M}$. Moreover, we say that $\left\langle X_{\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ is an $\mathcal{M}$-list of $\mathfrak{X}$ if $X_{\alpha}=\{m\in M:\mathcal{M}\models\psi(\alpha,m)\}$ for all $\alpha\in\mathrm{Ord}^{\mathcal{M}}.$
  2. $(ii)$
    Given an $\mathcal{M}$-listable $\mathfrak{X}\subseteq\mathbb{B}$, we say that $\mathfrak{X}$ is $\mathcal{M}$-thick if every $\mathcal{M}$-finite intersection of elements of $\mathfrak{X}$ is nonempty. More precisely, $\mathfrak{X}$ is $\mathcal{M}$-thick if $\mathcal{M}$ satisfies the following, where $\psi$ is a listing formula for $\mathfrak{X}$:
@@ 2. $(ii)$
$\forall k\in\omega$ $\forall\left\langle\alpha_{i}:i<k\right\rangle\in\mathrm{Ord}^{k}\ \exists\gamma\in\mathrm{Ord}\ \forall i<k\ \psi(\alpha_{i},\gamma).$
  1. $(iii)$
    For each $n\in\omega$, let $\mathbb{B}_{n}$ be the collection of $X\in\mathbb{B}$ such that $X$ is $\Sigma_{n}$-definable in $\mathcal{M}$ (parameters allowed). Note that by Facts 1 and 2, $\mathbb{B}_{n}$ is $\mathcal{M}$-listable for each $n\in\omega$.
  2. $(iv)$
[…]
We wish to construct an ultrafilter $\mathcal{U}$ $\subseteq$ $\mathbb{B}$ that satisfies the following properties:
  1. $(1)$
    $\mathcal{U}$ is an ultrafilter on $\mathbb{B}$.
  2. $(2)$
    $\mathcal{F\subseteq U}$.
  3. $(3)$
    Given any $\mathcal{M}$-definable $f:\mathrm{Ord}^{\mathcal{M}}\rightarrow\mathcal{M}$, there is some $P\in\mathcal{U}$ such that either $f\upharpoonright P$ is one-to-one, or the range of $f\upharpoonright P$ is a set (as opposed to a class).
  4. $(4)$
    Given any $n\in\mathbb{\omega}$, $\mathcal{U}\cap\mathbb{B}_{n}$ is $\mathcal{M}$-listable. This condition is equivalent to: for any $\mathcal{M}$-list $\left\langle X_{\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$, $\left\{\alpha\in\mathrm{Ord}^{\mathcal{M}}:X_{\alpha}\in\mathcal{U}\right\}\in\mathbb{B}.$
The desired ultrafilter $\mathcal{U}$ will be defined as $\bigcup\limits_{n\in\omega}\mathcal{U}_{n}$, where $\mathcal{U}_{n+1}$ will be defined from $\mathcal{U}_{n}$ using an internal recursion within $\mathcal{M}$ of length $\mathrm{Ord}^{\mathcal{M}}$. Thus, intuitively speaking, $\mathcal{U}$ will be constructed in $\omega\times\mathrm{Ord}^{\mathcal{M}}$-stages.
  • In what follows, for $n\in\omega$, we fix an $\mathcal{M}$-list $\left\langle f_{n,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ of all parametrically definable $\Sigma_{n}$-functions [^21] The the notion of $\mathcal{M}$-listability can be naturally extended to family $\mathcal{Y}$ of subsets of $M^{2}$, since $\mathcal{Y}$ can be coded as a family $\mathfrak{X}_{\mathcal{Y}}$ of subsets of $M$ with the help of a canonical pairing function. of $\mathcal{M}$ with domain $\mathrm{Ord}^{\mathcal{M}}$, and an $\mathcal{M}$-list $\left\langle S_{n,\alpha}:\alpha\in\mathrm{Ord}^{\mathcal{M}}\right\rangle$ of all elements of $\mathbb{B}_{n}$. Such $\mathcal{M}$-lists exist by Facts 1 and 2 listed at the beginning of Stage 1.
Let $\mathcal{U}_{0}$ $\mathcal{F}$. Thus $\mathcal{U}_{0}$ is $\mathcal{M}$-listable and $\mathcal{M}$-thick. We now describe the inductive construction of $\mathcal{U}_{n+1}$ from $\mathcal{U}_{n}$. So suppose $n\in\omega$, and we have $\mathcal{U}_{n}\subseteq\mathbb{B}$ that satisfies the clauses $\mathbb{C}_{1}(n)$ through $\mathbb{C}_{3}(n)$ below. Note that conditions $\mathbb{C}_{2}(n)$ and $\mathbb{C}_{3}(n)$ only kick in for $n\geq 1$ and ensure that $\mathcal{U}_{n}$ abides by certain $\Sigma_{n-1}$-obligations for $n\geq 1$.
  1. $\mathbb{C}_{1}(n):$
    $\mathcal{U}_{n}$ is $\mathcal{M}$-listable and $\mathcal{M}$-thick.
  2. $\mathbb{C}_{2}(n):$
    If $n\geq 1$, then $\forall\alpha\in\mathrm{Ord}^{\mathcal{M}}\ \exists X\in\mathcal{U}_{n}$ ($f_{n-1,\alpha}\upharpoonright X$ is one-to-one, or the range of $f_{n-1,\alpha}\upharpoonright X$ is a set).
  3. $\mathbb{C}_{3}(n):$
[…]
The following lemma is crucial for the construction of $\mathcal{U}_{n+1}\supseteq\mathcal{U}_{n}$ such that $\mathcal{U}_{n+1}$ satisfies $\mathbb{C}_{1}(n+1)$ through $\mathbb{C}_{3}(n+1).$
Lemma 3.1.1. Suppose $n\in\omega$, $\mathcal{F}\subseteq\mathfrak{X}\subseteq\mathbb{B},$ $\mathfrak{X}$ is $\Sigma_{n}$-listable in $\mathcal{M}$, and $\mathfrak{X}$ is $\mathcal{M}$-thick. Then there is a large enough $k\in\omega$ that satisfies the following two conditions:
  1. (a)
    Given any $A\subseteq\mathrm{Ord}^{\mathcal{M}}$ that is $\Sigma_{m}$-definable in $\mathcal{M}$, let $\widehat{A}\in\mathbb{B}$ be defined by: $\widehat{A}=A$ if $\mathfrak{X\cup\{}A\}$ is $\mathcal{M}$-thick, and otherwise $\A\widehat{A}:=\mathrm{Ord}^{\mathcal{M}}\backslash$ A. Then $\mathfrak{X\cup\{}\widehat{A}\}$ is $\mathcal{M}$-thick, and $\widehat{A}$ is $\Sigma_{p}$-definable in $\mathcal{M}$ for $p=\max\{m,n\}+k$.
  2. (b)
    Given any $f:\mathrm{Ord}^{\mathcal{M}}\rightarrow M$ that is $\Sigma_{m}$-definable in $\mathcal{M}$, there is some $P\in\mathbb{B}$ that is $\Sigma_{p}$-definable in $\mathcal{M}$ for $p=\max\{m,n\}+k$ such that $\mathfrak{X\cup\{}P\}$ is $\mathcal{M}$-thick, and $f\upharpoonright$ $P$ is either one-to-one, or the range of $f\upharpoonright$ $P$ is coded as an element of $\mathcal{M}$.
Proof. (a) is easy to see, so we leave the proof to the reader. Our proof of (b) will not explicitly specify the bound $k$, instead we focus on the combinatorics, which is uniform enough to yield the existence of such a bound. Given an $\mathcal{M}$-definable $f:\mathrm{Ord}^{\mathcal{M}}\rightarrow M$, we distinguish two cases:
Case 1: There is some $\gamma\in\mathrm{Ord}^{\mathcal{M}}$ such that $\mathfrak{X}\cup\{Y_{\gamma}\}$ is $\mathcal{M}$-thick, where
$Y_{\gamma}=\{\alpha\in\mathrm{Ord}^{\mathcal{M}}:\mathcal{M}\models f(\alpha)<\gamma\}.$